OpenStudy (anonymous):

After 6 years, an investment is worth $8479. If the money was invested at 9% interest compounded semiannually, what was the initial investment?

OpenStudy (anonymous):

\[P \left(\frac{R}{N}+1\right)^{\text{NT}}=\text{Amount} \text{Final}\]

OpenStudy (anonymous):

Final Amount=8479

OpenStudy (anonymous):

R=.09

OpenStudy (anonymous):

N=2

OpenStudy (anonymous):

T=6

OpenStudy (anonymous):

\[P \left(\frac{.09}{2}+1\right)^{2*6}=8479\]

OpenStudy (anonymous):

\[P =\frac{8479}{\left(\frac{.09}{2}+1\right)^{2*6}}\]

OpenStudy (anonymous):

P=4999.76